# DeTracker: Motion-decoupled Vehicle Detection and Tracking in Unstabilized Satellite Videos

### v1

@@ Abstract
Satellite videos provide continuous observations of surface dynamics but pose significant challenges for multi-object tracking (MOT), especially under unstabilized conditions where platform jitter and the weak appearance of tiny objects jointly degrade tracking performance. To address this problem, we propose DeTracker, a joint detection-and-tracking framework tailored for unstabilized satellite videos. DeTracker introduces a Global–Local Motion Decoupling (GLMD) module that explicitly separates satellite platform motion from true object motion through global alignment and local refinement, leading to improved trajectory stability and motion estimation accuracy. In addition, a Temporal Dependency Feature Pyramid (TDFP) module is developed to perform cross-frame temporal feature fusion, enhancing the continuity and discriminability of tiny-object representations. We further construct a new benchmark dataset, SDM-Car-SU, which simulates multi-directional and multi-speed platform motions to enable systematic evaluation of tracking robustness under varying motion perturbations. Extensive experiments on both simulated and real unstabilized satellite videos demonstrate that DeTracker significantly outperforms existing methods, achieving 61.1% MOTA on SDM-Car-SU and 47.3% MOTA on real satellite video data.
@@
Unstabilized satellite video, multi-object tracking, motion decoupling, temporal feature fusion.
@@ I Introduction
Object tracking in satellite videos is an important and emerging technique for understanding dynamic processes on the Earth’s surface. Compared with traditional static remote sensing imagery, satellite videos provide continuous spatio-temporal observations, which effectively support the long-term analysis of ground targets and demonstrate great potential in applications such as urban traffic monitoring [61], disaster emergency response [36], and smarty city management [44]. However, due to the high-altitude imaging nature of satellite platforms, targets in satellite videos are typically small in scale and densely distributed, posing significant challenges to multi-object tracking (MOT).
Early MOT approaches primarily follow the tracking-by-detection (TBD) paradigm [19, 9, 42, 41, 18], in which object detectors are used to localize targets in each frame, followed by cross-frame association based on re-identification features. With continued progress, spatio-temporal context modeling [56], cross-scale feature fusion [21, 4], and Transformer-based architectures [62, 57] have been incorporated to better handle complex background disturbances and capture long-range dependencies across frames. More recently, joint detection-and-tracking (JDT) frameworks have emerged as a dominant approach [16, 59, 31, 17]. They employ a unified backbone to predict object bounding boxes and identity embeddings across consecutive frames within a single forward pass, thereby performing both object localization and identity association jointly. By leveraging identity matching via embedding similarity, these methods reduce reliance on manually designed association rules and achieve more consistent MOT performance in large-scale dynamic scenes.
At present, most MOT methods for satellite videos are developed on pre-stabilized video sequences by video stablization methods [22, 49, 58], where background motion due to satellite platform movement is compensated to reduce its disturbances to object detection and improve the accuracy of object association. Despite these advances, most existing satellite video tracking methods still rely on a stabilize-then-track paradigm. In this paradigm, the stabilization stage typically involves computationally expensive motion estimation and resampling operations, which makes it primarily suitable for offline ground-based processing after data downlink and difficult to meet the requirements of current online target information extraction and service applications.
@@ I-1 Ambiguity in Decoupling of Mixed Motions
In unstabilized satellite videos, the observed motion arises from a superposition of true target motion and global motion induced by the satellite platform [55]. As illustrated in Fig. 1(a), even slight attitude perturbations can shift the sensor’s field of view, making originally static ground backgrounds appear highly dynamic. Moreover, the magnitude of platform-induced motion is typically larger than the displacement of tiny targets, thereby dominating the observed motion. This coupling between platform motion and object motion makes motion decoupling inherently ambiguous, which in turn degrade motion estimation and lead to tracking trajectory drift.
@@ I-2 Difficulty in Feature Extraction of Tiny Objects
To address the above challenges, we propose DeTracker, a joint detection-and-tracking framework tailored for unstabilized satellite videos. Unlike conventional approaches that rely on explicit image-level stabilization, DeTracker performs background motion compensation directly in the feature space, enabling the network to learn motion-aware representations even under strong global motion disturbances. Specifically, we design a Global-Local Motion Decoupling (GLMD) module that disentangles the global motion induced by the satellite platform from the true motion of the targets at the feature level, thereby suppressing background drift while preserving target-level motion cues critical for tracking. Within the GLMD module, a Global Alignment (GA) submodule leverages motion-oriented self-attention to capture long-range correspondences across consecutive frames, focusing on platform-induced background variations to estimate frame-level global transformations and achieve feature-level alignment. Complementing the global alignment, a Local Refinement (LR) submodule learns region-level motion offsets in the preliminarily aligned feature space, fine-tuning local feature misalignments caused by target motion and thereby enhancing the discriminability of tiny objects. Furthermore, we propose a Temporal Dependency Feature Pyramid (TDFP) module to improve feature extraction and modeling of tiny targets in unstabilized satellite videos, where frame-to-frame consistency is weakened. By performing multi-level temporal feature fusion and explicitly modeling cross-frame dependencies, TDFP enhances the spatio-temporal consistency of target representations, thereby improving detection robustness and tracking stability.
@@
    We propose DeTracker, a MOT framework for small targets in unstabilized satellite videos, aimed at on-orbit information services. Experimental results show that DeTracker achieves MOTA improvements of 5.1% and 8.3% on the simulated dataset and real satellite videos, respectively, outperforming existing state-of-the-art methods.
@@
    We introduce the Global–Local Motion Decoupling (GLMD) module and the Temporal Dependency Feature Pyramid (TDFP) module. The GLMD module decouples satellite platform motion from target motion via global alignment and local refinement, suppressing background drift, while the TDFP module enhances the temporal discriminability of small-target features through multi-scale aggregation and cross-frame dependency modeling.
@@
    To further advance research on tracking in unstabilized satellite videos, we construct a Simulated Unstabilized satellite video MOT benchmark, SDM-Car-SU, incorporating multi-directional and multi-velocity perturbations to emulate satellite platform motion. In addition, several real unstabilized satellite video clips are provided to evaluate algorithm performance in practical on-orbit scenarios.
@@ II-A Object Tracking in Generic Video
The JDT paradigm unify detection and association within a single network, generally yielding more stable identity consistency at the cost of increased training complexit and weaker cross-domain generalization. Early JDT methods leverage Siamese architectures to strengthen the discriminative representation of object features [16, 59], while subsequent works enhance identity consistency and tracking robustness by incorporating visibility modulation, two-stage association, and feature enhancement, thereby reducing ID switches [31, 17]. More recently, Transformer-based JDT methods introduce learnable identity embeddings and data-driven association mechanisms, replacing heuristic matching with end-to-end identity prediction [62, 57, 13].
@@ II-B Object Tracking in Remote Sensing Video
Research on object tracking in remote sensing video can be broadly divided into correlation filter–based methods and deep learning–based methods.
@@ II-B1 Correlation Filter-based Methods
Correlation filter–based methods are among the earliest techniques applied to UAV and satellite video tracking owing to their high computational efficiency. While they struggle under rapid and unstable UAV motion, the characteristics of satellite videos, such as tiny objects, low signal-to-noise ratios, complex backgrounds, and relatively slow motion, make correlation filters the dominant paradigm in the early and middle stages of satellite video tracking [19, 9]. Subsequent works enhance robustness by integrating multiple filters [15, 10], incorporating motion estimation and multi-feature modeling into the KCF framework for low-resolution imagery [51], or combining Kalman filtering with temporal consistency constraints to mitigate drift under occlusion and background clutter [34].
@@ II-B2 Deep Learning-based Methods
Benefiting from flexible viewpoints and wide coverage, UAV videos have become a major testbed for deep learning–based tracking. Siamese network–based methods significantly improve tracking accuracy and robustness [32, 33, 14], while scale-aware and motion-aware frameworks address object size variation and complex dynamics [6, 23, 63]. Transformer-based architectures further enhance modeling capacity by jointly improving computational efficiency and capturing long-range dependencies [2, 8, 43]. To cope with low-resolution inputs, recent works adopt knowledge distillation and feature compensation mechanisms to preserve performance while reducing computational cost [26, 48].
In satellite videos, objects are typically extremely small within the image, which makes feature extraction highly challenging and leads to severe feature ambiguity. To alleviate this, researchers have proposed end-to-end frameworks with enhanced feature extractors to better represent blurred tiny objects [42, 41, 18]. By incorporating cross-frame motion modeling and spatio-temporal fusion, these methods strengthen temporal continuity, while multi-scale cascaded enhancement modules further improve scale adaptability [38]. Other approaches exploit dynamic object masks to highlight tiny-object regions or employ sparse sampling strategies to correct feature relationships and suppress background interference [56]. Recent work adopts dual-frame input with heatmap-based output and lightweight matching to preserve identity consistency, thereby enhancing the discriminability of tiny objects under dynamic noise [21, 4]. Trajectory-aware tracking frameworks further exploit motion history to improve both accuracy and efficiency [5, 7].
In summary, although notable progress, satellite video tracking still pose substantial challenges arising from dense tiny objects, weak feature representations, and the strong coupling between platform motion and object trajectories, motivating the need for dedicated frameworks that explicitly decouple motion and enhance tiny-object representations.
@@ III SDM-Car-SU: An Unstabilized Satellite Video Vehicle Tracking Dataset
Figure 2: Visualization of inter-frame motion and annotation mapping under unstabilized conditions. The left columns present the frame-wise annotation mapping results under different types of unstabilized perturbations, while the rightmost column visualizes the simulated platform motion directions.
@@ III-A Dataset Characteristics
SDM-Car-SU targets vehicle tracking in unstabilized satellite videos, as shown in Fig. 2. By simulating different motion directions, velocities, and combined motion patterns, it more faithfully reflects the complex background motion encountered in on-orbit imaging, thereby posing greater challenges for both detection and tracking. Although DAF [47] considers simulated unstabilized data, it varies only background motion speed and ignores motion direction, and provides a single simulated sequence, which limits its applicability for training and benchmarking learning-based MOT models. In contrast, SDM-Car-SU offers a more comprehensive and diversified modeling of background motion along three dimensions:
@@ III-B Dataset Construction
The satellite data are sourced from the original SDM-Car dataset [54], which contains 99 high-quality video sequences acquired by the Luojia-3 01 satellite [39], with a spatial resolution of 0.75 m and an image size of $1920\times 1080$. By introducing inter-frame motion into the video frames to simulate satellite platform movement, we constructed the SDM-Car-SU dataset, which comprises three subsets with different motion rates: U1, U2, and U3, containing 292, 283, and 232 video clips, respectively. The train, validation, and test splits follow the same protocol as the SDM-Car dataset.
@@
Specifically, for motion parameter configuration, we follow the strategy of DAF [47] on the SkySat dataset [50], which has a spatial resolution of 1.0 m and background window speeds set to 0.5–2.0 pixels per frame. These values are chosen according to the average vehicle motion at that resolution, thereby covering background perturbation levels from slower to faster than typical vehicle motion. Following the same rationale, we statistically analyze the inter-frame pixel displacements of the same vehicle targets in SDM-Car [54] and obtain a mean motion speed of approximately 2 pixels per frame. Accordingly, we set the simulated background motion amplitudes to 1, 2, and 3 pixels per frame to create scenarios with different degrees of image instability. The construction process of the SDM-Car-SU dataset is detailed as follows:
@@ III-B3 Annotation Filtering and Mapping
To ensure the validity of the generated data, we retain only those sequences that satisfy the following criteria: (i) all cropped windows remain completely within the boundaries of the original image, and (ii) each frame contains at least one valid object annotation.
@@
where $(x_{n}^{\text{crop}},y_{n}^{\text{crop}})$ denotes the top-left coordinate of the cropping window, and $W_{\text{crop}}$, $H_{\text{crop}}$ are its width and height, respectively.
@@ IV Proposed Method
This section introduces the proposed DeTracker. We outline the tracking framework in Section IV-A. Then, Section IV-B presents the Global-Local Motion Decoupling (GLMD) module for motion disentanglement and Section IV-C describes the Temporal Dependency Feature Pyramid (TDFP) module for temporal feature enhancement. The objective functions are detailed in Section IV-D.

### v2

@@ Abstract
Satellite videos provide continuous observations of surface dynamics but pose significant challenges for multi-object tracking (MOT), especially under unstabilized conditions where platform jitter and the weak appearance of tiny objects jointly degrade tracking performance. To address this problem, we propose DeTracker, a joint-detection-and-tracking framework tailored for unstabilized satellite videos. DeTracker introduces a task-driven Global–Local Motion Decoupling (GLMD) module to address the motion imbalance between dominant platform motion and weak target motion. It suppresses background-dominated motion via global semantic alignment at the feature level and captures target-specific motion through local refinement, improving trajectory stability and identity consistency. In addition, a Temporal Dependency Feature Pyramid (TDFP) module is developed to perform cross-frame temporal feature fusion, enhancing the continuity and discriminability of tiny-object representations. We further construct a new benchmark dataset, SDM-Car-SU, which simulates multi-directional and multi-speed platform motions to enable systematic evaluation of tracking robustness under varying motion perturbations. Extensive experiments on both simulated and real unstabilized satellite videos demonstrate that DeTracker significantly outperforms existing methods, achieving 61.1% MOTA on SDM-Car-SU and 45.3% MOTA on real satellite video data. The code and dataset will be publicly available at https://github.com/alex-chenjiajun/DeTracker.
@@
Unstabilized satellite video, multi-object tracking (MOT), motion decoupling, temporal feature fusion.
@@ I Introduction
Object tracking in satellite videos is an important and emerging technique for understanding dynamic processes on the Earth’s surface  [25]. Compared with traditional static remote sensing imagery, satellite videos provide continuous spatio-temporal observations, which effectively support the long-term analysis of ground targets and demonstrate great potential in applications such as urban traffic monitoring [67], disaster emergency response [40], and smart city management [20]. However, due to the high-altitude imaging nature of satellite platforms, targets in satellite videos are typically small in scale and densely distributed, posing significant challenges to multi-object tracking (MOT).
Early MOT approaches primarily follow the tracking-by-detection (TBD) paradigm [19, 9, 49, 47, 18], in which object detectors are used to localize targets in each frame, followed by cross-frame association based on re-identification features. With continued progress, spatio-temporal context modeling [63, 22], cross-scale feature fusion [50, 4], and Transformer-based architectures [68, 64] have been incorporated to better handle complex background disturbances and capture long-range dependencies across frames. More recently, joint-detection-and-tracking (JDT) frameworks have emerged as a dominant approach [16, 65, 35, 17]. They employ a unified backbone to predict object bounding boxes and identity embeddings across consecutive frames within a single forward pass, thereby performing both object localization and identity association jointly. By leveraging identity matching via embedding similarity, these methods reduce reliance on manually designed association rules and achieve more consistent MOT performance in large-scale dynamic scenes.
At present, most MOT methods for satellite videos are developed on pre-stabilized video sequences by video stabilization methods [26, 56, 62], where background motion due to satellite platform movement is compensated to reduce its disturbances to object detection and improve the accuracy of object association. Despite these advances, most existing satellite video tracking methods still rely on a stabilize-then-track paradigm. In this paradigm, the stabilization stage typically involves computationally expensive motion estimation and resampling operations, which makes it primarily suitable for offline ground-based processing after data downlink and difficult to meet the requirements of current online target information extraction and service applications.
@@ I-1 Ambiguity in Decoupling of Mixed Motions
In unstabilized satellite videos, the observed motion arises from a superposition of true target motion and global motion induced by the satellite platform [62]. As illustrated in Fig. 1(a), even slight attitude perturbations can shift the sensor’s field of view, making originally static ground backgrounds appear highly dynamic. Moreover, the magnitude of platform-induced motion is typically larger than the displacement of tiny targets, thereby dominating the observed motion. This coupling between platform motion and object motion makes motion decoupling inherently ambiguous, which in turn degrades motion estimation and leads to tracking trajectory drift.
@@ I-2 Difficulty in Feature Extraction of Tiny Objects
To address the above challenges, we propose DeTracker, a joint-detection-and-tracking framework tailored for unstabilized satellite videos. Unlike conventional approaches that rely on explicit image-level stabilization, DeTracker performs background motion compensation directly in the feature space, enabling the network to learn motion-aware representations even under strong global motion disturbances. Specifically, to address the motion coupling issue, we design a Global–Local Motion Decoupling (GLMD) module that explicitly decouples the global motion induced by the satellite platform from the target’s intrinsic motion at the feature level, thereby suppressing background drift while preserving target-level motion cues critical for tracking. Within the GLMD module, a Global Alignment (GA) submodule leverages a motion-oriented self-attention mechanism to model long-range semantic dependencies across consecutive frames, focusing on suppressing platform-induced background feature variations to achieve cross-frame semantic-level alignment. Complementing the global alignment, a Local Refinement (LR) submodule learns region-level motion offsets in the preliminarily aligned feature space, capturing the fine-grained target motion information that was originally masked by background drift and thereby enhancing the discriminability of tiny objects. Furthermore, we propose a Temporal Dependency Feature Pyramid (TDFP) module to improve feature extraction and modeling of tiny targets in unstabilized satellite videos, where frame-to-frame consistency is weakened. By performing multi-level temporal feature fusion and explicitly modeling cross-frame dependencies, TDFP enhances the spatio-temporal consistency of target representations, thereby improving detection robustness and tracking stability.
@@
    We propose DeTracker, an MOT framework for small targets in unstabilized satellite videos, tailored for on-orbit information services. Experimental results show that DeTracker achieves MOTA improvements of 5.1% and 8.3% on the simulated dataset and real satellite videos, respectively, outperforming existing state-of-the-art methods.
@@
    We propose the Global–Local Motion Decoupling (GLMD) module and the Temporal Dependency Feature Pyramid (TDFP) module. The GLMD explicitly decouples platform motion from target motion via feature-level global alignment and local refinement, effectively suppressing background drift for robust tracking. The TDFP models cross-frame dependencies and aggregates multi-scale features to enhance the temporal discriminability of small targets.
@@
    To facilitate research on tracking in unstabilized satellite videos, we construct a Simulated Unstabilized satellite video MOT benchmark, SDM-Car-SU, incorporating multi-directional and multi-velocity perturbations to emulate satellite platform motion. In addition, several real unstabilized satellite video clips are provided to evaluate algorithm performance in practical on-orbit scenarios.
@@ II-A Object Tracking in Generic Video
The JDT paradigm unifies detection and association within a single network, typically achieving more stable identity consistency at the cost of higher training complexity and weaker cross-domain generalization. Early JDT methods leverage Siamese architectures to strengthen the discriminative representation of object features [16, 65], while subsequent works enhance identity consistency and tracking robustness by incorporating visibility modulation, two-stage association, and feature enhancement, thereby reducing ID switches [35, 17]. More recently, Transformer-based JDT methods introduce learnable identity embeddings and data-driven association mechanisms, replacing heuristic matching with end-to-end identity prediction [68, 64, 13].
@@ II-B Object Tracking in Remote Sensing Video
Research on object tracking in remote sensing video can be broadly divided into correlation filter-based methods and deep learning-based methods.
@@ II-B1 Correlation Filter-based Methods
Correlation filter-based methods are among the earliest techniques applied to UAV and satellite video tracking owing to their high computational efficiency. While they struggle under rapid and unstable UAV motion, the characteristics of satellite videos, such as tiny objects, low signal-to-noise ratios, complex backgrounds, and relatively slow motion, make correlation filters the dominant paradigm in the early and middle stages of satellite video tracking [19, 9]. Subsequent works enhance robustness by integrating multiple filters [15, 10], incorporating motion estimation and multi-feature modeling into the KCF framework for low-resolution imagery [58], or combining Kalman filtering with temporal consistency constraints to mitigate drift under occlusion and background clutter [38].
@@ II-B2 Deep Learning-based Methods
Benefiting from flexible viewpoints and wide coverage, UAV videos have become a major testbed for deep learning-based tracking. Siamese network-based methods significantly improve tracking accuracy and robustness [36, 37, 14], while scale-aware and motion-aware frameworks address object size variation and complex dynamics [23, 27, 69]. Transformer-based architectures further enhance modeling capacity by jointly improving computational efficiency and capturing long-range dependencies [2, 8, 51]. To cope with low-resolution inputs, recent works adopt knowledge distillation and feature compensation mechanisms to preserve performance while reducing computational cost [30, 55]. In addition, several studies have explored spatio-temporal feature interaction mechanisms to alleviate frequent ID switches caused by small object scales, sparse feature representations, and significant appearance variations in UAV scenarios. By introducing spatio-temporal trajectory priors to enhance feature association, these methods effectively mitigate visual feature degradation and improve identity consistency [46, 45].
In satellite videos, objects are typically extremely small within the image, which makes feature extraction highly challenging and leads to severe feature ambiguity. To alleviate this, researchers have proposed end-to-end frameworks with enhanced feature extractors to better represent blurred tiny objects [49, 47, 18]. By incorporating cross-frame motion modeling and spatio-temporal fusion, these methods strengthen temporal continuity, while multi-scale cascaded enhancement modules further improve scale adaptability [42]. To further advance motion modeling, a motion-inspired cross-modality learning mechanism has been introduced [6]. By mining global motion flows to capture latent motion patterns, this approach bridges the semantic gap between motion and visual features, significantly improving the representation of dynamic tiny objects. Other approaches exploit dynamic object masks to highlight tiny-object regions or employ sparse sampling strategies to correct feature relationships and suppress background interference [63]. Building on the concept of sparsity, recent methods utilizing sparse spatio-temporal point cloud representations reduce redundant background clutter and enhance the detection of tiny, low-contrast objects [48]. Recent works adopt dual-frame input with heatmap-based output and lightweight matching to preserve identity consistency, thereby enhancing the discriminability of tiny objects under dynamic noise [24, 4]. Trajectory-aware tracking frameworks further exploit motion history to improve both accuracy and efficiency [5, 7].
In summary, although notable progress, satellite video tracking still poses substantial challenges arising from dense tiny objects, weak feature representations, and the strong coupling between platform motion and object trajectories, motivating the need for dedicated frameworks that explicitly decouple motion and enhance tiny-object representations.
@@ III SDM-Car-SU: An Unstabilized Satellite Video Vehicle Tracking Dataset
Figure 2: Visualization of inter-frame motion, annotation mapping, and target trajectories under unstabilized conditions. The left columns present the frame-wise annotation mapping results along with the coupled motion trajectories over the preceding 20 frames under different types of unstabilized perturbations, while the rightmost column visualizes the simulated platform motion directions.
@@ III-A Dataset Characteristics
SDM-Car-SU targets vehicle tracking in unstabilized satellite videos, as shown in Fig. 2. By simulating different motion directions, velocities, and combined motion patterns, it more faithfully reflects the complex background motion encountered in on-orbit imaging, thereby posing greater challenges for both detection and tracking. Although DAF [54] considers simulated unstabilized data, it varies only background motion speed, neglects motion direction, and provides just one simulated sequence, which limits its applicability for training and benchmarking learning-based MOT models. In contrast, SDM-Car-SU offers a more comprehensive and diversified modeling of background motion along three dimensions:
@@ III-B Dataset Construction
The satellite data are sourced from the original SDM-Car dataset [61], which contains 99 high-quality video sequences acquired by the Luojia-3 01 satellite [43], with a spatial resolution of 0.75 m and an image size of $1920\times 1080$. By introducing inter-frame motion into the video frames to simulate satellite platform movement, we construct the SDM-Car-SU dataset, which comprises three subsets with different motion rates: U1, U2, and U3, containing 335, 265, and 220 video clips, respectively. The train, validation, and test splits follow the same protocol as the SDM-Car dataset.
@@
Specifically, for motion parameter configuration, we follow the strategy of DAF [54] on the SkySat dataset [57], which has a spatial resolution of 1.0 m and background window speeds set to 0.5–2.0 pixels per frame. These values are chosen according to the average vehicle motion at that resolution, thereby covering background perturbation levels from slower to faster than typical vehicle motion. Following the same rationale, we statistically analyze the inter-frame pixel displacements of the same vehicle targets in SDM-Car [61] and obtain a mean motion speed of approximately 2 pixels per frame. Accordingly, we set the simulated background motion amplitudes to 1, 2, and 3 pixels per frame to create scenarios with different degrees of image instability. Statistical analysis of 211 unstabilized satellite video clips from the Luojia-3 01 satellite validates the configured motion amplitudes. As demonstrated in Fig. 3, approximately 90% of the inter-frame motion magnitudes fall within a 3-pixel range. This empirical distribution provides robust substantiation for the chosen simulation parameters. The construction process of the SDM-Car-SU dataset is detailed as follows:
Figure 3: Distribution of inter-frame motion amplitudes in real-world unstabilized videos from the Luojia-3 01 satellite.
@@ III-B3 Annotation Filtering and Mapping
To ensure the validity of the generated data, we retain only those sequences that satisfy the following criteria: (i) all cropped windows remain completely within the boundaries of the original image; and (ii) each frame contains at least one valid object annotation.
@@
where $(x_{n}^{\text{crop}},y_{n}^{\text{crop}})$ is the top-left coordinate of the cropping window, and $W_{\text{crop}}$ and $H_{\text{crop}}$ denote its width and height.
@@ IV Proposed Method
This section introduces the proposed DeTracker. We outline the tracking framework in Section IV-A. Then, Section IV-B presents the Global–Local Motion Decoupling (GLMD) module for motion disentanglement and Section IV-C describes the Temporal Dependency Feature Pyramid (TDFP) module for temporal feature enhancement. The objective functions are detailed in Section IV-D.